# Stable non-Hausdorff Riemannian foliations

### v1

@@ 1. Introduction
A foliation $\mathcal{F}$ on a compact manifold $M$ is called stable if any other foliation sufficiently close to $\mathcal{F}$ in the $C^{\infty}$-topology is conjugate to $\mathcal{F}$ under a diffeomorphism of $M$. Here the $\mathcal{C}^{\infty}$-topology on the space of $k$-dimensional foliations is induced by the inclusion
@@
Recent work [4] by the author and F. Zeiser aimed to find examples of stable foliations other than those considered by Hamilton. To explain the main result of [4], recall that for any foliation $\mathcal{F}$ there is a canonical representation $\nabla$ of the Lie algebroid $T\mathcal{F}$ on the normal bundle $N\mathcal{F}$, called the Bott connection. Hence, one obtains a complex $\big{(}\Omega^{\bullet}(\mathcal{F},N\mathcal{F}),d_{\nabla}\big{)}$ with cohomology groups $H^{\bullet}(\mathcal{F},N\mathcal{F})$. Heitsch [6] showed that first order deformations of the foliation $\mathcal{F}$, modulo those obtained by applying an isotopy to $\mathcal{F}$, are given by elements in the cohomology group $H^{1}(\mathcal{F},N\mathcal{F})$. Therefore, vanishing of $H^{1}(\mathcal{F},N\mathcal{F})$ can be interpreted heuristically as the infinitesimal requirement for stability of $\mathcal{F}$. The main result in [4] states that the vanishing of $H^{1}(\mathcal{F},N\mathcal{F})$ actually implies that $\mathcal{F}$ is stable, at least when $\mathcal{F}$ is Riemannian. Recall that $\mathcal{F}$ is Riemannian if $M$ admits a metric that is bundle-like with respect to $\mathcal{F}$ [13]. Equivalently, $\mathcal{F}$ is locally defined by Riemannian submersions.
@@ 2. Suspension foliations
the associated suspension $(B_{\varphi},\mathcal{F})$ is a Riemannian foliation (in fact, a Lie $G$-foliation). We will only consider homomorphisms $\varphi$ for which the image $H:=\varphi(\pi_{1}(B))$ is a dense subgroup of $G$. In that case, the suspension foliation $\mathcal{F}$ on $B_{\varphi}$ is a Lie $G$-foliation with dense leaves. This follows immediately from the fact that $\mathcal{F}$ is a transversely parallelizable foliation on a compact connected manifold, all of whose basic functions are constant [10, Thm. 4.24]. To see this, note that a frame of left invariant vector fields on $G$ gives a transverse parallelism for the horizontal foliation on $\widehat{B}\times G$, which descends to a transverse parallelism for $\mathcal{F}$ on $B_{\varphi}$. Moreover, the $\mathcal{F}$-basic functions on $B_{\varphi}$ are constant since
@@ 3.2. Actions of Kazhdan groups
Kazhdan’s property (T) is relevant for our purpose because of the following: a countable discrete group $\Gamma$ has property (T) exactly when the first group cohomology $H^{1}(\Gamma,V)$ vanishes for every orthogonal representation of $\Gamma$ on a real Hilbert space $V$ (see [1, Chapter 2]). This statement is not directly applicable in our situation because the representations that we encounter are no Hilbert spaces, see (2). To overcome this issue, we need to invoke the following result by Lubotzky-Zimmer [8, Theorem 4.1]. We only state a particular case of their result, which holds under more general assumptions that are implied by property (T).
@@ 3.3. Proof of Main Theorem
The action of $\pi_{1}(B)$ on $G$ is isometric with respect to any left invariant Riemannian metric on $G$. Moreover, it is well-known that if $\Gamma\subset G$ is a dense subgroup of a compact Lie group, then the action by left translations $\Gamma\curvearrowright G$ is ergodic with respect to the normalized Haar measure on $G$ [12, Lemma 2.4.1]. Since by assumption $\varphi(\pi_{1}(B))\subset G$ is dense, it follows that the action $\pi_{1}(B)\curvearrowright(G,\mu)$ is ergodic. Moreover, the fundamental group $\pi_{1}(B)$ of a compact manifold $B$ is finitely generated. Applying Prop. 3.7, we see that the cohomology group (3) vanishes. Stability of $\mathcal{F}$ immediately follows from Thm. 1.2 in the Introduction. ∎
@@ 4. An example
The aim of this section is to construct a discrete group $\Gamma$ which can serve as the fundamental group $\pi_{1}(B)$ in Thm. 3.8. We will obtain $\Gamma$ as a suitable arithmetic lattice in $SO(3,2)$ which can be realized as a dense subgroup of $SO(5)$. The lattice in question was considered before in connection with the Ruziewicz problem, which concerns rotationally invariant measures on spheres [7, §3.4].
@@ 4.1. Arithmetic lattices
We start by recalling an arithmetic construction of lattices, following [9, Chapter IX, §1.7]. Fix an integer $n\geq 3$ and a subfield $K\subset\mathbb{C}$ which is a finite extension of $\mathbb{Q}$. Denote by $d=[K:\mathbb{Q}]$ the degree of the extension. Assume that
@@
[9, Chapter IX, §1.7] Assume that $\mathcal{T}\neq\mathcal{R}$ and that the group $SO_{\Phi}$ is almost $K$-simple. Then $SO_{\Phi}(\mathcal{O})$ is an arithmetic lattice in $\prod_{\sigma\in\mathcal{S}}SO_{\Phi_{\sigma}}(k_{\sigma})$.
@@ 4.2. An example
  • Lattices in connected semisimple Lie groups are finitely presented [3, §1.1]. Viewing $\Gamma$ as a lattice inside $SO(3,2)$, this implies that $\Gamma$ is finitely presented, even though $SO(3,2)$ has $2$ connected components. To see this, let us denote by $SO(3,2)^{+}$ its identity component, which is a normal index $2$ subgroup $SO(3,2)^{+}\subset SO(3,2)$. The fact that $SO(3,2)^{+}$ is an open subgroup implies that $\Gamma\cap SO(3,2)^{+}$ is a lattice inside $SO(3,2)^{+}$, hence $\Gamma\cap SO(3,2)^{+}$ is finitely presented by the above. Moreover, since
@@
  • We have the compact, connected Lie group $G:=SO(5)$ and a group homomorphism $\varphi:\Gamma\rightarrow SO(5)$ given by the restriction of the second projection in $SO(3,2)\times SO(5)$. It was shown in [7, Prop. 3.4.3] that the projection of $\Gamma$ in $SO(5)$ is dense.
@@
We obtain an arithmetic lattice $\Gamma\subset SO(n-2,2)\times SO(n)$ which projects to a lattice in $SO(n-2,2)$ and a dense subgroup in $SO(n)$ [7, Prop. 3.4.3]. Moreover, $\Gamma$ has property (T).
The method fails for $n<5$. In fact, Thm. 3.8 admits no examples in which $G=SO(n)$ with $n<5$. Indeed, if $\varphi:\Gamma\rightarrow SO(n)$ is a group homomorphism where $\Gamma$ is a discrete Kazhdan group and $n<5$, then the image $\varphi(\Gamma)$ is finite. This was shown by Zimmer in [14, Thm. 7] for $SO(3)$ and $SO(4)$. It is immediate in the case of $SO(2)$, because of the following. Since $SO(2)$ is abelian, the group homomorphism $\varphi:\Gamma\rightarrow SO(2)$ factors through the abelianization Γ/[$\Gamma/[\Gamma,\Gamma]$. Because the abelianization of a discrete group with property (T) is finite [1, Cor. 1.3.6], it follows that the image $\varphi(\Gamma)\subset SO(2)$ is finite. This shows that the suspension foliation of any homomorphism $\varphi:\Gamma\rightarrow SO(n)$, with $\Gamma$ a discrete Kazhdan group and $n<5$, is Hausdorff since all of its leaves are compact (see §2).

### v2

@@ 1. Introduction
A foliation $\mathcal{F}$ on a compact manifold $M$ is called stable if any other foliation sufficiently close to $\mathcal{F}$ in the $C^{\infty}$-topology is conjugate to $\mathcal{F}$ under a diffeomorphism of $M$. Here the $C^{\infty}$-topology on the space of $k$-dimensional foliations is induced by the inclusion
@@
In fact, various versions of this result have been obtained independently throughout the years, using a variety of techniques. Langevin and Rosenberg [9] established $C^{1}$-stability of fibrations whose fiber $L$ satisfies $H^{1}(L)=0$. Epstein and Rosenberg [4] proved the analog of Thm. 1.1 for Hausdorff $C^{k}$-foliations with compact leaves on manifolds that are not necessarily compact. Finally, Del Hoyo and Fernandes [2] proved the statement of Thm. 1.1 for smooth paths of foliations $\mathcal{F}_{t}$ deforming $\mathcal{F}$, rather than foliations $\mathcal{F}^{\prime}$ close to $\mathcal{F}$.
Recent work by the author and Zeiser [6] aimed to find examples of stable foliations other than the Hausdorff foliations considered above. To explain the main result of [6], recall that for any foliation $\mathcal{F}$ there is a canonical representation $\nabla$ of the Lie algebroid $T\mathcal{F}$ on the normal bundle $N\mathcal{F}$, called Bott connection. Hence, one obtains a complex $\big{(}\Omega^{\bullet}(\mathcal{F},N\mathcal{F}),d_{\nabla}\big{)}$ with cohomology groups $H^{\bullet}(\mathcal{F},N\mathcal{F})$. Heitsch [8] showed that first order deformations of the foliation $\mathcal{F}$, modulo those obtained by applying an isotopy to $\mathcal{F}$, are given by elements in the cohomology group $H^{1}(\mathcal{F},N\mathcal{F})$. Therefore, vanishing of $H^{1}(\mathcal{F},N\mathcal{F})$ can be interpreted heuristically as the infinitesimal requirement for stability of $\mathcal{F}$. The main result in [6] states that the vanishing of $H^{1}(\mathcal{F},N\mathcal{F})$ actually implies that $\mathcal{F}$ is stable, at least when $\mathcal{F}$ is Riemannian. Recall that $\mathcal{F}$ is Riemannian if $M$ admits a metric that is bundle-like with respect to $\mathcal{F}$ [16]. Equivalently, $\mathcal{F}$ is locally defined by Riemannian submersions.
@@ 2. Suspension foliations
the associated suspension $(B_{\varphi},\mathcal{F})$ is a Riemannian foliation (in fact, a Lie $G$-foliation). We will only consider homomorphisms $\varphi$ for which the image $H:=\varphi(\pi_{1}(B))$ is a dense subgroup of $G$. In that case, the suspension foliation $\mathcal{F}$ on $B_{\varphi}$ is a Lie $G$-foliation with dense leaves. This follows immediately from the fact that $\mathcal{F}$ is a transversely parallelizable foliation on a compact connected manifold, all of whose basic functions are constant [13, Thm. 4.24]. To see this, note that a frame of left invariant vector fields on $G$ gives a transverse parallelism for the horizontal foliation on $\widehat{B}\times G$, which descends to a transverse parallelism for $\mathcal{F}$ on $B_{\varphi}$. Moreover, the $\mathcal{F}$-basic functions on $B_{\varphi}$ are constant since
@@ 3.2. Actions of Kazhdan groups
Kazhdan’s property (T) is relevant for our purpose because of the following: a countable discrete group $\Gamma$ has property (T) exactly when the first group cohomology $H^{1}(\Gamma,V)$ vanishes for every orthogonal representation of $\Gamma$ on a real Hilbert space $V$ (see [1, Chapter 2]). This statement is not directly applicable in our situation because the representations that we encounter are no Hilbert spaces, see (2). To overcome this issue, we need to invoke the following result by Lubotzky-Zimmer [11, Theorem 4.1]. We only state a particular case of their result, which holds under more general assumptions that are implied by property (T).
@@ 3.3. Proof of Main Theorem
The action of $\pi_{1}(B)$ on $G$ is isometric with respect to any left invariant Riemannian metric on $G$. Moreover, it is well-known that if $\Gamma\subset G$ is a dense subgroup of a compact Lie group, then the action by left translations $\Gamma\curvearrowright G$ is ergodic with respect to the normalized Haar measure on $G$ [15, Lemma 2.4.1]. Since by assumption $\varphi(\pi_{1}(B))\subset G$ is dense, it follows that the action $\pi_{1}(B)\curvearrowright(G,\mu)$ is ergodic. Moreover, the fundamental group $\pi_{1}(B)$ of a compact manifold $B$ is finitely generated. Applying Prop. 3.7, we see that the cohomology group (3) vanishes. Stability of $\mathcal{F}$ immediately follows from Thm. 1.2 in the Introduction. ∎
@@ 4. An example
The aim of this section is to construct a discrete group $\Gamma$ which can serve as the fundamental group $\pi_{1}(B)$ in Thm. 3.8. We will obtain $\Gamma$ as a suitable arithmetic lattice in $SO(3,2)$ which can be realized as a dense subgroup of $SO(5)$. The lattice in question was considered before in connection with the Ruziewicz problem, which concerns rotationally invariant measures on spheres [10, §3.4].
@@ 4.1. Arithmetic lattices
We start by recalling an arithmetic construction of lattices, following [12, Chapter IX, §1.7]. Fix an integer $n\geq 3$ and a subfield $K\subset\mathbb{C}$ which is a finite extension of $\mathbb{Q}$. Denote by $d=[K:\mathbb{Q}]$ the degree of the extension. Assume that
@@
[12, Chapter IX, §1.7] Assume that $\mathcal{T}\neq\mathcal{R}$ and that the group $SO_{\Phi}$ is almost $K$-simple. Then $SO_{\Phi}(\mathcal{O})$ is an arithmetic lattice in $\prod_{\sigma\in\mathcal{S}}SO_{\Phi_{\sigma}}(k_{\sigma})$.
@@ 4.2. An example
  • Lattices in connected semisimple Lie groups are finitely presented [5, §1.1]. Viewing $\Gamma$ as a lattice inside $SO(3,2)$, this implies that $\Gamma$ is finitely presented, even though $SO(3,2)$ has $2$ connected components. To see this, let us denote by $SO(3,2)^{+}$ its identity component, which is a normal index $2$ subgroup $SO(3,2)^{+}\subset SO(3,2)$. The fact that $SO(3,2)^{+}$ is an open subgroup implies that $\Gamma\cap SO(3,2)^{+}$ is a lattice inside $SO(3,2)^{+}$, hence $\Gamma\cap SO(3,2)^{+}$ is finitely presented by the above. Moreover, since
@@
  • We have the compact, connected Lie group $G:=SO(5)$ and a group homomorphism $\varphi:\Gamma\rightarrow SO(5)$ given by the restriction of the second projection in $SO(3,2)\times SO(5)$. It was shown in [10, Prop. 3.4.3] that the projection of $\Gamma$ in $SO(5)$ is dense.
@@
We obtain an arithmetic lattice $\Gamma\subset SO(n-2,2)\times SO(n)$ with property (T), which projects to a dense subgroup in $SO(n)$ [10, Prop. 3.4.3].
The method fails for $n<5$. In fact, Thm. 3.8 admits no examples in which $G=SO(n)$ with $n<5$. Indeed, if $\varphi:\Gamma\rightarrow SO(n)$ is a group homomorphism where $\Gamma$ is a discrete Kazhdan group and $n<5$, then the image $\varphi(\Gamma)$ is finite. This was shown by Zimmer in [17, Thm. 7] for $SO(3)$ and $SO(4)$. It is immediate in the case of $SO(2)$, because of the following. Since $SO(2)$ is abelian, the group homomorphism $\varphi:\Gamma\rightarrow SO(2)$ factors through the abelianization Γ/[$\Gamma/[\Gamma,\Gamma]$. Because the abelianization of a discrete group with property (T) is finite [1, Cor. 1.3.6], it follows that the image $\varphi(\Gamma)\subset SO(2)$ is finite. This shows that the suspension foliation of any homomorphism $\varphi:\Gamma\rightarrow SO(n)$, with $\Gamma$ a discrete Kazhdan group and $n<5$, is Hausdorff since all of its leaves are compact (see §2).